# Finite energy well-posedness for nonlinear Schrödinger equations with non-vanishing conditions at infinity

### v3

@@ 3.1. Local well-posedness
for all $n\in{\mathbf{N}}$, we obtain from (3.16) that there exists $\phi\in L^{\infty}([0,T];L^{r_{1}}({\mathbf{R}}^{2}))$ such that $|\psi_{q,n}|\leq\phi$ a.e. on $[0,T_{2})\times{\mathbf{R}}^{2}$. Therefore,
$\left|\left(G_{\infty}(\psi^{\ast})-G_{\infty}(\psi_{n})\right)\right|\,|\nabla\psi^{\ast}|\leq C|\nabla\psi^{\ast}|\in L^{1}([0,T);L^{2}({\mathbf{R}}^{2})),$
$\left|\left(G_{q}(\psi^{\ast})-G_{q}(\psi_{n})\right)\right|\,|\nabla\psi^{\ast}|\leq C\left(|\psi^{\ast}|^{2\alpha}+|\phi|^{2\alpha}\right)|\nabla\psi^{\ast}|\in L^{q_{1}^{\prime}}([0,T);L^{r_{1}^{\prime}}({\mathbf{R}}^{2})),$
so that the dominated convergence Theorem then implies that (3.27) is violated. The inequality (3.17) follows for the time interval $[0,T_{2}]$ where we stress that $T_{2}>0$ only depends on $M$.
@@
×$\times\left(d_{\mathbb{E}}(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}^{1},\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}^{\ast})+\|\Phi(\psi_{1})-\Phi(\psi^{\ast})\|_{L_{t}^{\infty}H_{x}^{1}}\right)$
@@
Step 4: Note that Step 3 yields continuous dependence on the initial data w.r.t. to the topology of $\mathbb{E}$ induced by the metric $d_{\mathbb{E}}$ on a time interval $[0,T_{2}]$ where $T_{2}$ only depends on $M$. One may hence cover $[0,T]$ by the union of intervals $[t_{k},t_{k+1}]$ with $t_{k}=kT_{2}$ for $k\in\{0,...,N-1\}$ with $N=\lceil\frac{T}{T_{2}}\rceil$ finite. For all $\varepsilon>0$, there exists $\delta_{N}>0$ such that $d_{\mathbb{E}}(\psi_{1}(t_{N-1}),\psi^{\ast}(t_{N-1}))<\delta_{N}$ yields $\sup_{t\in[t_{N-1},T]}d_{\mathbb{E}}(\psi_{1}(t),\psi^{\ast}(t))<\varepsilon$. Next, there exists $\delta_{N-1}>0$ such that $d_{\mathbb{E}}(\psi_{1}(t_{N-2}),\psi^{\ast}(t_{N-2}))<\delta_{N-1}$ yields $\sup_{t\in[t_{N-2},t_{N-1}]}d_{\mathbb{E}}(\psi_{1}(t),\psi^{\ast}(t))<\delta_{N}$. One may then iterate the scheme finitely many times in order to recover $\delta=\delta_{1}>0$ such that $d_{\mathbb{E}}(\psi_{1}^{0},\psi_{0}^{\ast})<\delta$ implies $\sup_{t\in[0,T]}d_{\mathbb{E}}(\psi_{1}(t),\psi^{\ast}(t))<\varepsilon$.
@@
$\left\|G_{1}(\psi)\partial_{t}\psi+G_{2}(\psi)\overline{\partial_{t}\psi}\right\|_{N^{0}([0,T]\times{\mathbf{R}}^{2})}\leq C\|G_{\infty}(\psi)|\partial_{t}\psi|\|_{L_{t}^{1}L_{x}^{2}}+\|G_{q}(\psi)|\partial_{t}\psi|\|_{N^{0}}\\ \leq C\|\partial_{t}\psi\|_{L_{t}^{1}L_{x}^{2}}+\|\left(1+|\psi|^{2\alpha}\right)|\partial_{t}\psi|\|_{N^{0}}\leq CT\|\partial_{t}\psi\|_{L_{t}^{\infty}L_{x}^{2}}+T^{\frac{1}{q_{1}^{\prime}}}Z_{T}(\psi)^{2\alpha}\|\partial_{t}\psi\|_{L_{t}^{\infty}L_{x}^{2}}.$
@@ 3.2. Global well-posedness
Note that to complete the proof of the theorem it suffices to prove that the Hamiltonian energy is conserved for all solutions $\psi\in C([0,T^{\ast});\mathbb{E}({\mathbf{R}}^{2}))$. Global existence then follows by arguing as in the proof of Corollary 3.6. To that end, given initial data $\psi_{0}\in\mathbb{E}({\mathbf{R}}^{3})$ and the unique solution $\psi\in C([0,T^{\ast});\mathbb{E}({\mathbf{R}}^{2}))$ to (1.1) such that $\psi(0)=\psi_{0}$, we observe that thanks to Lemma 2.10 there exists $\{\psi_{0}^{n}\}\subset\mathbb{E}({\mathbf{R}}^{2})\cap C^{\infty}({\mathbf{R}}^{2})$ such that $\Delta\psi_{0}^{n}\in L^{2}({\mathbf{R}}^{2})$ and $d_{\mathbb{E}}(\psi_{0},\psi_{0}^{n})$ converges to 00 as $n$ goes to infinity. Lemma 3.5 provides a sequence of unique global solutions $\psi_{n}\in C({\mathbf{R}},\mathbb{E}({\mathbf{R}}^{2}))$ such that $\mathcal{H}(\psi_{n})(t)=\mathcal{H}(\psi_{0}^{n})$ for all $n$. Relying on the continuous dependence on the initial data, we conclude that for any $0<T<T^{\ast}$ one has
@@
Let $d=3$ and $f$ be such that Assumption 1.1 is satisfied. Then,
@@
$\left\|\mathcal{N}_{2,\infty}(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u)\right\|_{L_{t}^{1}L_{x}^{2}}\leq CT\mathcal{L}^{3}\left(\operatorname{supp}(1-\eta(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u))\right)^{\frac{1}{2}}\leq CT\left(2\sqrt{2R}+M\right)$
@@
$\left\|\mathcal{N}_{2,q}(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u)\right\|_{L_{t}^{1}L_{x}^{2}+L_{t}^{q^{\prime}}L_{x}^{r^{\prime}}}\\ \leq\left\|(1+|\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u|^{2\alpha})|\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u|(1-\chi(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u))\right\|_{L_{t}^{1}L_{x}^{2}+L_{t}^{q^{\prime}}L_{x}^{r^{\prime}}}\\ \leq CT(2\sqrt{2R}+M)+\left\|\left|\left(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u\right)(1-\chi(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u))\right|^{2\alpha+1}\right\|_{L_{t}^{q^{\prime}}L_{x}^{r^{\prime}}}\\ \leq CT(2\sqrt{2R}+M)+C{T}^{\frac{q-q^{\prime}}{qq^{\prime}}}\left\|(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u)_{q}\right\|_{L^{\infty}L^{r}}^{2\alpha}\left\|(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u)_{q}\right\|_{L_{t}^{q}L_{x}^{r}}\\ \leq C\left(T+T^{\frac{q-q^{\prime}}{qq^{\prime}}}\left(2\sqrt{2R}+M\right)^{2\alpha}\right)\left(2\sqrt{2R}+M\right).$
Moreover, Assumption 1.1, see also (2.33), imply the bound
@@
$\left\|\nabla\mathcal{N}_{1}(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u)+\nabla\mathcal{N}_{2,\infty}(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u)\right\|_{L_{t}^{1}L_{x}^{2}}\\ \leq CT\left(\|\nabla\psi_{0}\|_{L_{t}^{\infty}L_{x}^{2}}+\|\nabla u\|_{L_{t}^{\infty}L_{x}^{2}}\right)\leq CT\left(2\sqrt{2R}+M\right).$
To control $\nabla\mathcal{N}_{2,q}$, note that $\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\nabla\psi_{0}\in L^{q}([0,T];L^{r}({\mathbf{R}}^{3}))$ for any admissible pair $(q,r)$ from Lemma 2.14 and $\mathcal{E}(\psi_{0})\leq R$. Therefore,
$\|\nabla\mathcal{N}_{2,q}(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u)\|_{L_{t}^{1}L_{x}^{2}+L_{t}^{q^{\prime}}L_{x}^{r^{\prime}}}\leq CT\left(\|\nabla\psi_{0}\|_{L^{2}}+\|u\|_{X_{T}}\right)\\ +C\left(\left\||(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u)_{q}|^{2\alpha}\nabla\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}\right\|_{L_{t}^{q^{\prime}}L_{x}^{r^{\prime}}}+\left\||(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u)_{q}|^{2\alpha}\nabla u\right\|_{L_{t}^{q^{\prime}}L_{x}^{r^{\prime}}}\right)\\ \leq CT(2\sqrt{2R}+M)+CT^{\frac{q-q^{\prime}}{qq^{\prime}}}(2\sqrt{2R}+M)^{2\alpha}\left(\|\nabla\psi_{0}\|_{L_{x}^{2}}+\|\nabla u\|_{L_{t}^{q}L_{x}^{r}}\right)\\ \leq C\left(T+T^{\frac{q-q^{\prime}}{qq^{\prime}}}(2\sqrt{2R}+M)^{2\alpha}\right)\left(2\sqrt{2R}+M\right).$
@@
$\left\|\mathcal{N}_{2,\infty}\left(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u\right)-\mathcal{N}_{2,\infty}\left(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+v\right)\right\|_{L_{t}^{1}L_{x}^{2}}\leq CTd_{X}(u,v).$
@@
$\left\|\mathcal{N}_{2,q}\left(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u\right)-\mathcal{N}_{2,q}\left(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+v\right)\right\|_{L^{1}L^{2}+L_{t}^{r^{\prime}}L_{x}^{q^{\prime}}}$
@@
Therefore, it suffices to set $M=\sqrt{R}$ and to choose $T=T(M)>0$ sufficiently small in order to conclude that $\Phi:X_{T}\rightarrow X_{T}$ and $\Phi$ defines a contraction on $X_{T}$. The Banach fixed-point Theorem yields a unique solution $u\in X_{T}$ to (4.2). In particular, $\psi(t)=\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u(t)$ solves (1.1) with $\psi\in C([0,T];\mathbb{E}({\mathbf{R}}^{3}))$.
Uniqueness For $R>0$ fixed, let $\psi_{0}\in\mathbb{E}({\mathbf{R}}^{3})$ with $\mathcal{E}(\psi_{0})\leq R$ and $\psi_{1},\psi_{2}\in C([0,T];\mathbb{E}({\mathbf{R}}^{3}))$ two solutions to (1.1) such that $\psi_{1}(0)=\psi_{2}(0)=\psi_{0}$. We note that $\psi_{1}-\psi_{2}\in S^{1}([0,T]\times{\mathbf{R}}^{3})$. In particular, from the Strichartz estimate (2.25) and arguing as for the local existence we obtain that
@@
Thus, there exists $T_{1}>0$ sufficiently small such that $\psi_{1}=\psi_{2}$ a.e. on $[0,T_{1}]\times{\mathbf{R}}^{3}$. As $T_{1}$ only depends on $Z_{T}(\psi_{i})$ with $i=1,2$ one may iterate the argument. This yields uniqueness in $C([0,T];\mathbb{E}({\mathbf{R}}^{3}))$.
Blow up alternative The proof of the blow-up alternative follows verbatim the proof of the respective statement for $d=2$, see Proposition 3.2 and is omitted.
@@
Given $\psi_{0}\in\mathbb{E}({\mathbf{R}}^{3})$, let $\psi\in C([0,T^{\ast});\mathbb{E}({\mathbf{R}}^{3}))$ be the unique maximal solution to (1.1) with initial data $\psi_{0}$. Then, there exists $|c|=1$ and $v\in C([0,T^{\ast});\mathcal{F}_{c})$ such that $\psi(t)=c+v(t)$ for all $t\in[0,T^{\ast})$ and where $v$ is solution to
@@
Note that Lemma 4.2 reduces the well-posedness of (1.1) in $\mathbb{E}({\mathbf{R}}^{3})$ to solving the affine problem (4.3) in $\mathcal{F}_{c}$ where the constant $c$ is determined by the choice of the initial data. In particular, the continuous dependence on the initial data can be stated equivalently in terms of the metric (2.7) with the constants $c_{1},c_{2}$ determined by the initial data.
If the nonlinearity is such that $f$ satisfies (1.14), then it is convenient to implement the well-posedness result in homogeneous spaces by exploiting Strichartz estimates on the gradient, see also [23, Remark 4.5] for (1.6) and [36, Proposition 1.1.18] for (1.1) with nonlinearity (1.4). Indeed, Assumption (1.14) ensures that $\nabla\mathcal{N}$ is locally Lipschitz. A suitable choice of the functional spaces for the local well-posedness is given by
@@
where the Strichartz admissible pair is for instance $(q,r)=(10,\frac{30}{13})$, see [36, Proposition 1.1.18].
However, in the framework of Assumption 1.1, this is ruled out by the lack of regularity of the nonlinearity $f$. More precisely, for $\nabla\mathcal{N}$ to be locally Lipschitz we require (1.14).
@@
Let $f$ satisfy Assumption 1.1, $T>0$, $(q,r)$ as defined in (4.1) and $\psi_{1},\psi_{2}\in C([0,T];\mathbb{E}({\mathbf{R}}^{3}))$ such that $\psi_{i}=c_{i}+v_{i}$ with $c_{i}\in{\mathbf{C}}$, $|c_{i}|=1$ and $v_{i}\in C([0,T];\mathcal{F}_{c})$ for $i=1,2$. Then, there exists $\theta\in(0,1]$ such that
@@
$\|\mathcal{N}_{2,\infty}(\psi_{1})-\mathcal{N}_{2,\infty}(\psi_{2})\|_{L_{t}^{1}L_{x}^{2}}\leq CT\left(Z_{T}(\psi_{1})+Z_{T}(\psi_{2})\right)|c_{1}-c_{2}|\\ +CT^{\frac{1}{2}}\left(Z_{T}(\psi_{1})^{\frac{2}{3}}+{Z_{T}(\psi_{2})}^{\frac{2}{3}}\right)\left\|v_{1}-v_{2}\right\|_{L_{t}^{2}L_{x}^{6}}.$
Third, we show the desired bound for $\mathcal{N}_{2,q}(\psi_{1})-\mathcal{N}_{2,q}(\psi_{2})$. As $|\psi_{i}|\geq\frac{3}{2}$ on $\operatorname{supp}(\mathcal{N}_{2,q}(\psi_{i}))$, it follows from (2.36) that
$\left|\mathcal{N}_{2,q}(\psi_{1})-\mathcal{N}_{2,q}(\psi_{2})\right|\leq C\left(1+|\psi_{1}|^{2\alpha}+|\psi_{2}|^{2\alpha}\right)|\psi_{1}-\psi_{2}|$
@@
$\left|\mathcal{N}_{2,q}(\psi_{1})-\mathcal{N}_{2,q}(\psi_{2})\right|\leq C\left(1+|\psi_{1,q}|^{\alpha}+|\psi_{2,q}|^{\alpha}\right)|\psi_{1}-\psi_{2}|,$

### v4

@@ 3.1. Local well-posedness
for all $n\in{\mathbf{N}}$, we obtain from (3.16) that there exists $\phi\in L^{\infty}([0,T];L^{r_{1}}({\mathbf{R}}^{2}))$ such that $|\psi_{q,n}|\leq\phi$ a.e. on $0,T_{2})\times{\mathbf{R}}^{2}$. Therefore,
$\left|\left(G_{\mathrm{bd}}(\psi^{\ast})-G_{\mathrm{bd}}(\psi_{n})\right)\right|\,|\nabla\psi^{\ast}|\leq C|\nabla\psi^{\ast}|\in L^{1}([0,T);L^{2}({\mathbf{R}}^{2})),$
$\left|\left(G_{\mathrm{int}}(\psi^{\ast})-G_{\mathrm{int}}(\psi_{n})\right)\right|\,|\nabla\psi^{\ast}|\leq C\left(|\psi^{\ast}|^{2\alpha}+|\phi|^{2\alpha}\right)|\nabla\psi^{\ast}|\in L^{q_{1}^{\prime}}([0,T);L^{r_{1}^{\prime}}({\mathbf{R}}^{2})),$
so that the dominated convergence Theorem implies that (3.27) is violated. The inequality (3.17) follows for the time interval $0,T_{2}]$ where we stress that $T_{2}>0$ only depends on $M$.
@@
$\times\left(d_{\mathbb{E}}(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}^{1},\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}^{\ast})+\|\Phi(\psi_{1})-\Phi(\psi^{\ast})\|_{L_{t}^{\infty}H_{x}^{1}}\right)$
@@
Step 4: Note that Step 3 yields continuous dependence on the initial data w.r.t. to the topology of $\mathbb{E}$ induced by the metric $d_{\mathbb{E}}$ on a time interval $0,T_{2}]$ where $T_{2}$ only depends on $M$. One may hence cover $[0,T]$ by the union of intervals $t_{k},t_{k+1}]$ with $t_{k}=kT_{2}$ for $k\in\{0,...,N-1\}$ with $N=\lceil\frac{T}{T_{2}}\rceil$ finite. For all $\varepsilon>0$, there exists $\delta_{N}>0$ such that $d_{\mathbb{E}}(\psi_{1}(t_{N-1}),\psi^{\ast}(t_{N-1}))<\delta_{N}$ yields $\sup_{t\in[t_{N-1},T]}d_{\mathbb{E}}(\psi_{1}(t),\psi^{\ast}(t))<\varepsilon$. Next, there exists $\delta_{N-1}>0$ such that $d_{\mathbb{E}}(\psi_{1}(t_{N-2}),\psi^{\ast}(t_{N-2}))<\delta_{N-1}$ yields $\sup_{t\in[t_{N-2},t_{N-1}]}d_{\mathbb{E}}(\psi_{1}(t),\psi^{\ast}(t))<\delta_{N}$. One may then iterate the scheme finitely many times in order to recover the existence of a $\delta=\delta_{1}>0$ such that $d_{\mathbb{E}}(\psi_{1}^{0},\psi_{0}^{\ast})<\delta$ implies $\sup_{t\in[0,T]}d_{\mathbb{E}}(\psi_{1}(t),\psi^{\ast}(t))<\varepsilon$.
@@
$\left\|G_{1}(\psi)\partial_{t}\psi+G_{2}(\psi)\overline{\partial_{t}\psi}\right\|_{N^{0}([0,T]\times{\mathbf{R}}^{2})}\leq C\|G_{\mathrm{bd}}(\psi)|\partial_{t}\psi|\|_{L_{t}^{1}L_{x}^{2}}+\|G_{\mathrm{int}}(\psi)|\partial_{t}\psi|\|_{N^{0}}\\ \leq C\|\partial_{t}\psi\|_{L_{t}^{1}L_{x}^{2}}+\|\left(1+|\psi|^{2\alpha}\right)|\partial_{t}\psi|\|_{N^{0}}\leq CT\|\partial_{t}\psi\|_{L_{t}^{\infty}L_{x}^{2}}+T^{\frac{1}{q_{1}^{\prime}}}Z_{T}(\psi)^{2\alpha}\|\partial_{t}\psi\|_{L_{t}^{\infty}L_{x}^{2}}.$
@@ 3.2. Global well-posedness
Note that to complete the proof of the theorem it suffices to prove that the Hamiltonian energy is conserved for all solutions $\psi\in C([0,T^{\ast});\mathbb{E}({\mathbf{R}}^{2}))$. Global existence then follows by arguing as in the proof of Corollary 3.6. To that end, given initial data $\psi_{0}\in\mathbb{E}({\mathbf{R}}^{3})$ and the unique solution $\psi\in C([0,T^{\ast});\mathbb{E}({\mathbf{R}}^{2}))$ to (1.1) such that $\psi(0)=\psi_{0}$, we observe that thanks to Lemma 2.10 there exists $\{\psi_{0}^{n}\}\subset\mathbb{E}({\mathbf{R}}^{2})\cap C^{\infty}({\mathbf{R}}^{2})$ such that $\Delta\psi_{0}^{n}\in L^{2}({\mathbf{R}}^{2})$ and $d_{\mathbb{E}}(\psi_{0},\psi_{0}^{n})$ converges to 0 as $n$ goes to infinity. Lemma 3.5 provides a sequence of unique global solutions $\psi_{n}\in C({\mathbf{R}},\mathbb{E}({\mathbf{R}}^{2}))$ such that $\mathcal{H}(\psi_{n})(t)=\mathcal{H}(\psi_{0}^{n})$ for all $n$. Relying on the continuous dependence on the initial data, we conclude that for any $0<T<T^{\ast}$ one has
@@
Let $d=3$ and $f$ be as in Assumption 1.1.
@@
$\left\|\mathcal{N}_{2,\mathrm{bd}}(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u)\right\|_{L_{t}^{1}L_{x}^{2}}\leq CT\mathcal{L}^{3}\left(\operatorname{supp}(1-\eta(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u))\right)^{\frac{1}{2}}\leq CT\left(2\sqrt{2R}+M\right)$
@@
$\left\|\mathcal{N}_{2,\mathrm{int}}(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u)\right\|_{L_{t}^{1}L_{x}^{2}+L_{t}^{q^{\prime}}L_{x}^{r^{\prime}}}\\ \leq\left\|(1+|\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u|^{2\alpha})|\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u|(1-\chi(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u))\right\|_{L_{t}^{1}L_{x}^{2}+L_{t}^{q^{\prime}}L_{x}^{r^{\prime}}}\\ \leq CT(2\sqrt{2R}+M)+\left\|\left|\left(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u\right)(1-\chi(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u))\right|^{2\alpha+1}\right\|_{L_{t}^{q^{\prime}}L_{x}^{r^{\prime}}}\\ \leq CT(2\sqrt{2R}+M)+C{T}^{\frac{q-q^{\prime}}{qq^{\prime}}}\left\|(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u)_{q}\right\|_{L^{\infty}L^{r}}^{2\alpha}\left\|(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u)_{q}\right\|_{L_{t}^{q}L_{x}^{r}}\\ \leq C\left(T+T^{\frac{q-q^{\prime}}{qq^{\prime}}}\left(2\sqrt{2R}+M\right)^{2\alpha}\right)\left(2\sqrt{2R}+M\right).$
Moreover, Assumption 1.1, see also (2.32), implies the bound
@@
$\left\|\nabla\mathcal{N}_{1}(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u)+\nabla\mathcal{N}_{2,\mathrm{bd}}(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u)\right\|_{L_{t}^{1}L_{x}^{2}}\\ \leq CT\left(\|\nabla\psi_{0}\|_{L_{t}^{\infty}L_{x}^{2}}+\|\nabla u\|_{L_{t}^{\infty}L_{x}^{2}}\right)\leq CT\left(2\sqrt{2R}+M\right).$
To control $\nabla\mathcal{N}_{2,\mathrm{int}}$, note that $\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\nabla\psi_{0}\in L^{q}([0,T];L^{r}({\mathbf{R}}^{3}))$ for any admissible pair $(q,r)$ from Lemma 2.14 and $\mathcal{E}(\psi_{0})\leq R$. Therefore,
$\|\nabla\mathcal{N}_{2,\mathrm{int}}(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u)\|_{L_{t}^{1}L_{x}^{2}+L_{t}^{q^{\prime}}L_{x}^{r^{\prime}}}\leq CT\left(\|\nabla\psi_{0}\|_{L^{2}}+\|u\|_{X_{T}}\right)\\ +C\left(\left\||(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u)_{q}|^{2\alpha}\nabla\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}\right\|_{L_{t}^{q^{\prime}}L_{x}^{r^{\prime}}}+\left\||(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u)_{q}|^{2\alpha}\nabla u\right\|_{L_{t}^{q^{\prime}}L_{x}^{r^{\prime}}}\right)\\ \leq CT(2\sqrt{2R}+M)+CT^{\frac{q-q^{\prime}}{qq^{\prime}}}(2\sqrt{2R}+M)^{2\alpha}\left(\|\nabla\psi_{0}\|_{L_{x}^{2}}+\|\nabla u\|_{L_{t}^{q}L_{x}^{r}}\right)\\ \leq C\left(T+T^{\frac{q-q^{\prime}}{qq^{\prime}}}(2\sqrt{2R}+M)^{2\alpha}\right)\left(2\sqrt{2R}+M\right).$
@@
$\left\|\mathcal{N}_{2,\mathrm{bd}}\left(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u\right)-\mathcal{N}_{2,\mathrm{bd}}\left(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+v\right)\right\|_{L_{t}^{1}L_{x}^{2}}\leq CTd_{X}(u,v).$
@@
$\left\|\mathcal{N}_{2,\mathrm{int}}\left(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u\right)-\mathcal{N}_{2,\mathrm{int}}\left(\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+v\right)\right\|_{L^{1}L^{2}+L_{t}^{r^{\prime}}L_{x}^{q^{\prime}}}$
@@
Therefore, it suffices to set $M=\sqrt{R}$ and choose $T=T(M)>0$ sufficiently small in order to conclude that $\Phi:X_{T}\rightarrow X_{T}$ and $\Phi$ defines a contraction on $X_{T}$. The Banach fixed-point Theorem yields a unique solution $u\in X_{T}$ to (4.2). In particular, $\psi(t)=\mathrm{e}^{\frac{\mathrm{i}}{2}t\Delta}\psi_{0}+u(t)$ solves (1.1) with $\psi\in C([0,T];\mathbb{E}({\mathbf{R}}^{3}))$.
Uniqueness Let $R>0$ be fixed. Let $\psi_{0}\in\mathbb{E}({\mathbf{R}}^{3})$ with $\mathcal{E}(\psi_{0})\leq R$ and $\psi_{1},\psi_{2}\in C([0,T];\mathbb{E}({\mathbf{R}}^{3}))$ two solutions to (1.1) such that $\psi_{1}(0)=\psi_{2}(0)=\psi_{0}$. We note that $\psi_{1}-\psi_{2}\in S^{1}([0,T]\times{\mathbf{R}}^{3})$. In particular, from the Strichartz estimate (2.24) and arguing as for the local existence we obtain that
@@
Thus, there exists a sufficiently small $T_{1}>0$ such that $\psi_{1}=\psi_{2}$ a.e. on $0,T_{1}]\times{\mathbf{R}}^{3}$. As $T_{1}$ only depends on $Z_{T}(\psi_{i})$ with $i=1,2$ one may iterate the argument. This yields uniqueness in $C([0,T];\mathbb{E}({\mathbf{R}}^{3}))$.
Blow up alternative The proof of the blow-up alternative follows verbatim the proof of the respective statement for $d=2$, see Proposition 3.2 and thus it is omitted.
@@
Given $\psi_{0}\in\mathbb{E}({\mathbf{R}}^{3})$, let $\psi\in C([0,T^{\ast});\mathbb{E}({\mathbf{R}}^{3}))$ be the unique maximal solution to (1.1) with initial data $\psi_{0}$. Then, there exist $|c|=1$ and $v\in C([0,T^{\ast});\mathcal{F}_{c})$ such that $\psi(t)=c+v(t)$ for all $t\in[0,T^{\ast})$ and where $v$ is a solution to
@@
Note that Lemma 4.2 reduces the well-posedness of (1.1) in $\mathbb{E}({\mathbf{R}}^{3})$ to solving the affine problem (4.3) in $\mathcal{F}_{c}$ where the constant $c$ is determined by the choice of the initial data. In particular, the continuous dependence on the initial data can be stated equivalently in terms of the metric (2.9) with the constants $c_{1}$ and $c_{2}$ determined by the initial data.
If the nonlinearity is such that $f$ satisfies (1.14), then it is convenient to implement the well-posedness result in homogeneous spaces by exploiting Strichartz estimates on the gradient, see also [26, Remark 4.5] for (1.6) and [39, Proposition 1.1.18] for (1.1) with nonlinearity (1.4). Indeed, Assumption (1.14) ensures that $\nabla\mathcal{N}$ is locally Lipschitz. A suitable choice of the functional spaces for the local well-posedness is given by
@@
where the Strichartz admissible pair is for instance $(q,r)=(10,\frac{30}{13})$, see [39, Proposition 1.1.18].
However, note that in the framework of Assumption 1.1, this is ruled out by the lack of regularity of the nonlinearity $f$. More precisely, for $\nabla\mathcal{N}$ to be locally Lipschitz we require (1.14).
@@
Let $f$ satisfy Assumption 1.1, $T>0$, $(q,r)$ as defined in (4.1) and $\psi_{1},\psi_{2}\in C([0,T];\mathbb{E}({\mathbf{R}}^{3}))$ such that $\psi_{i}=c_{i}+v_{i}$ with $c_{i}\in{\mathbf{C}}$, $|c_{i}|=1$ and $v_{i}\in C([0,T];\mathcal{F}_{c_{i}})$ for $i=1,2$. Then, there exists $\theta\in(0,1]$ such that
@@
$\|\mathcal{N}_{2,\mathrm{bd}}(\psi_{1})-\mathcal{N}_{2,\mathrm{bd}}(\psi_{2})\|_{L_{t}^{1}L_{x}^{2}}\leq CT\left(Z_{T}(\psi_{1})+Z_{T}(\psi_{2})\right)|c_{1}-c_{2}|\\ +CT^{\frac{1}{2}}\left(Z_{T}(\psi_{1})^{\frac{2}{3}}+{Z_{T}(\psi_{2})}^{\frac{2}{3}}\right)\left\|v_{1}-v_{2}\right\|_{L_{t}^{2}L_{x}^{6}}.$
Third, we establish the desired bound for $\mathcal{N}_{2,\mathrm{int}}(\psi_{1})-\mathcal{N}_{2,\mathrm{int}}(\psi_{2})$. As $|\psi_{i}|\geq\frac{3}{2}$ on $\operatorname{supp}(\mathcal{N}_{2,\mathrm{int}}(\psi_{i}))$, it follows from (2.35) that
$\left|\mathcal{N}_{2,\mathrm{int}}(\psi_{1})-\mathcal{N}_{2,\mathrm{int}}(\psi_{2})\right|\leq C\left(1+|\psi_{1}|^{2\alpha}+|\psi_{2}|^{2\alpha}\right)|\psi_{1}-\psi_{2}|$
@@
$\left|\mathcal{N}_{2,\mathrm{int}}(\psi_{1})-\mathcal{N}_{2,\mathrm{int}}(\psi_{2})\right|\leq C\left(1+|\psi_{1,\mathrm{int}}|^{\alpha}+|\psi_{2,\mathrm{int}}|^{\alpha}\right)|\psi_{1}-\psi_{2}|,$